OpenStudy (anonymous):

Write and equation of a circle with the given canter a radius. Center:(3,4) Radius:6

OpenStudy (sidsiddhartha):

hint: equation of a circle with centre at (a,b) and having radius c is \[(x-a)^{2}+(y-b)^{2}=c ^{2}\]

OpenStudy (anonymous):

So it would be (x-3)^2 + (y-4)^2=6

OpenStudy (sidsiddhartha):

(x-3)^2 + (y-4)^2=6^2=36

OpenStudy (anonymous):

Thank You!